OpenStudy (anonymous):

how to solve this without a calculator sqrt (2009*2011*2015*2017+36)+10

OpenStudy (nurali):

Welcome to Openstudy.

OpenStudy (anonymous):

\[\sqrt{2009*2011*2015*2017+36}+10\] how to describe simply

OpenStudy (nurali):

\[\sqrt{2009*2011*2015*2017+36}+10\] \[4052159+10\] \[4052169\]

OpenStudy (skullpatrol):

@Nurali all done without a calculator and a single word of $$\Huge \text{EXPLANATION}$$

OpenStudy (anonymous):

Write it using Latex ?

OpenStudy (skullpatrol):

Write what?

OpenStudy (anonymous):

Prime Factorization Break all numbers down into prime factors, then match them. \(sqrt[100]=10 \times 10 = 2 \times 5 \times 2 \times 5\) There are two 2's and two 5's, therefore the answer is \(2\times 5 = 10\)

OpenStudy (asnaseer):

hint: try to first rewrite 2009, 2011, 2015 and 2017 in terms of 2013, e.g. 2009=(2013-4)

OpenStudy (anonymous):

OK I think expression under the root is a square.\[a=\frac{ 2009+2011+2015+2017 }{ 4 }=2013\] 2009⋅2011⋅2015⋅2017=(a−4)(a−2)(a+2)(a+4) thanks.

OpenStudy (anonymous):

rest is simple: \[a=2013^{2}=4052169\]